OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x

OpenStudy (anonymous):

@xapproachesinfinity if you're not busy, could you help me with this one? It is my last question!

OpenStudy (anonymous):

@ganeshie8 do you think you could help?

OpenStudy (xapproachesinfinity):

hmm do you know double angle identities? you should know them

OpenStudy (xapproachesinfinity):

sin2x=2sinxcosx

OpenStudy (xapproachesinfinity):

also cos^2 x= cos^2x-sin^2x

OpenStudy (xapproachesinfinity):

try to use those...

OpenStudy (xapproachesinfinity):

by the way is that \(\sin 2x -\cos2x\) or \(\sin^2 x-\cos^2 x\)

OpenStudy (xapproachesinfinity):

?